# DynamicVL: Benchmarking Multimodal Large Language Models for Dynamic City Understanding

### v1

@@ 6 Conclusion
In this paper, we presented DVL-Suite, a comprehensive framework for analyzing long-term urban dynamics through remote sensing imagery. Our dataset uniquely features 15,063 high-resolution multi-temporal images spanning 42 U.S. megacities, organized with detailed annotations across seven urban understanding tasks. This rich collection enables systematic evaluation of both pixel-precise change detection and comprehensive temporal understanding capabilities. Through extensive experiments with 15 state-of-the-art models, we identified significant challenges in long-term temporal understanding and quantitative analysis. To address these limitations, we introduced DVL-Instruct and developed DVLChat as a baseline model, demonstrating the potential of our framework for advancing dynamic city understanding.
Acknowledgement
This work was supported in part by the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), KAKENHI (25K03145). This work was also supported by NVIDIA Academic Grant. This work used computational resources Miyabi supercomputer provided by The University of Tokyo through Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures and High Performance Computing Infrastructure in Japan (Project ID: jh250017). Weihao Xuan is supported by RIKEN Junior Research Associate (JRA) Program.

### v2

@@ 7 Conclusion
In this paper, we present DVL-Suite, a large-scale vision-language benchmark for analyzing long-term urban dynamics through remote sensing imagery. Featuring 14,871 high-resolution multi-temporal images across 42 U.S. cities with detailed annotations spanning six urban understanding tasks, DVL-Suite enables systematic evaluation of MLLMs’ capabilities from pixel-precise change detection to comprehensive temporal reasoning. Through extensive evaluation of 18 state-of-the-art models, we reveal critical insights: current MLLMs struggle significantly with long-term temporal understanding and quantitative analysis, while scaling model parameters alone proves insufficient without domain-specific training data. To address these limitations, we introduce DVL-Instruct, a specialized instruction-tuning dataset, and develop DVLChat as a baseline model that demonstrates substantial improvements, showcasing the potential of domain-specific data for advancing multi-temporal urban understanding capabilities.
Acknowledgements
This work was supported in part by the Council for Science, Technology and Innovation (CSTI) and the Cross-ministerial Strategic Innovation Promotion Program (SIP) “Development of a Resilient Smart Network System against Natural Disasters” (funding agency: NIED), KAKENHI (25K03145). This work was also supported by NVIDIA Academic Grant. This work used computational resources on the Miyabi supercomputer, provided by The University of Tokyo through the Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures and High Performance Computing Infrastructure in Japan (Project ID: jh250017). Weihao Xuan is supported by RIKEN Junior Research Associate (JRA) Program.